# Mimicking Human Intuition: Cognitive Belief-Driven Reinforcement Learning

### v3

@@ 6.1 Empirical Evaluations in Discrete Action Space
Figure 2: Learning curves of CBDDQN, DQN, DDQN, PPO and Duel_DQN. First row for Box2D (CarRacing and LunarLander) Classic Control(CartPole and Acrbot). Second and Third row for Metadrive with 8 different maps.
To evaluate the effectiveness of value-based reinforcement learning under cognitive abstraction, we introduce CBDQ as a belief-augmented extension of Q-learning through modifying DQN algorithm. We benchmark CBDDQN against standard baselines—DQN, DDQN, Dueling DQN, and PPO—across representative environments from Classic Control and Box2D. As shown in Figure 2, CBDDQN consistently demonstrates superior sample efficiency and final performance. In tasks such as CartPole, Acrobot, and CarRacing, CBDDQN attains convergence within the first 20% of training iterations while maintaining the highest observed reward levels throughout learning. To evaluate generalization and adaptability in complex domains, we test CBDDQN on the MetaDrive benchmark, which includes diverse driving scenarios with high-dimensional inputs and dynamic environments. CBDDQN consistently outperforms all baselines, showing stable learning and higher cumulative rewards. Unlike PPO, which often converges suboptimally or shows instability, CBDDQN achieves steady improvement and superior final performance across all settings. Supplementary experiments under varying traffic densities and accident probabilities further confirm CBDQ’s robustness in high-risk decision-making settings (Appendix E).
To assess the impact of conceptual belief integration in on-policy reinforcement learning, we compare CBDPPO and standard PPO across eight Atari environments. As shown in Figure 3, CBDPPO consistently achieves higher initial returns and exhibits smoother training dynamics. In tasks such as AirRaid, Breakout, and Asteroids, performance increases rapidly within the first 5–10 million steps and subsequently stabilizes with low variance. In contrast, PPO often exhibits continued fluctuation or degradation in later training stages. The elevated starting performance observed in CBDPPO is attributed to the incorporation of concept-level priors into the initial policy. Specifically, a set of belief distributions is constructed from early episodes using latent clustering, yielding action preferences associated with high-reward states. These priors are integrated into the policy via a convex combination at each timestep, biasing exploration toward semantically coherent behaviors from the outset. Additionally, the belief integration weight $\beta_{t}$ increases linearly gradually during training, allowing the agent to transition from task-specific learning to concept-guided refinement. This, combined with the clipped surrogate objective, effectively regularizes policy updates. Once the blended policy enters the trust region, gradient steps become smaller, resulting in reduced variance and stable convergence. These results suggest that conceptually structured priors can improve both the sample efficiency and stability of on-policy optimization, supporting the practical utility of the CBD-RL framework in large-scale and complex control settings.
6.2 Empirical Evaluations in Continuous Action Space
Table 1 presents experimental results across four carefully selected continuous action space environments. These environments represent a spectrum of human-like motor control tasks: from precise end-effector manipulation (Reacher), to coordinated limb movement (Bipedal Walker), to complex multi-joint locomotion (Ant and Humanoid). The results demonstrate CBDSAC’s superior performance in capturing human-like control strategies. Particularly in Ant and Humanoid environments, which demand sophisticated balance and coordination similar to natural movement patterns, CBDSAC maintains consistently higher rewards throughout training. While SAC achieves comparable final performance in Reacher and Bipedal Walker, CBDSAC exhibits notably faster convergence, suggesting more efficient learning of natural movement primitives. This accelerated learning aligns with human motor learning patterns, where basic movement principles are quickly adapted to specific tasks.
Figure 3: Learning curves of CBDPPO and PPO on selected Atari tasks. CBDPPO exhibits higher initial performance and lower variance in later stages, indicating improved sample efficiency and training stability through concept-guided policy regularization.
7 Research Insight
Effective experience utilization is central to reinforcement learning—not merely as training data, but as a mechanism for bridging past behavior with future decisions. This shift from reactive reward maximization to structured inference echoes Sutton’s view of experience as the foundation of intelligent systems (Dohare et al., 2024; Silver & Sutton, 2025), and aligns with cognitive science findings that humans generalize and infer from sparse data with remarkable efficiency (Tenenbaum et al., 2011; Lake et al., 2015; Tenenbaum et al., 2006). CBD-RL embodies this perspective through cognitively inspired mechanisms, yet remains an approximation of human cognition. Current belief–reward integration relies on a fixed convex combination (Eq. 8, 18), whereas human inference is context-sensitive and often modeled by Bayesian model averaging or hierarchical inference (Griffiths & Tenenbaum, 2007; Ma et al., 2006). Incorporating adaptive fusion—e.g., precision-weighted updates or meta-learned gating—could improve alignment with human-like inference under uncertainty. Similarly, the CCF module employs Euclidean clustering, which overlooks the non-Euclidean structure of human conceptual spaces shaped by semantic, attentional, and causal factors (Nosofsky, 1986; Nickel & Kiela, 2017; Murphy, 2004). Enhancing CCF with learned similarity metrics—such as contrastive objectives, hyperbolic embeddings, or causal constraints—offers a promising path toward bridging cognitive theory and representation learning in RL.
@@ 8 Conclusion
The Cognitive Belief-Driven Reinforcement Learning framework advances RL by incorporating cognitive-inspired mechanisms. By leveraging conceptual clustering to organize experiences efficiently and employing probabilistic reasoning for decision-making under uncertainty, CBD-RL improves sample efficiency while maintaining computational scalability. Its implementations, CBDQ, CBDPPO and CBDSAC, excel in both discrete and continuous action spaces. This framework bridges cognitive science and reinforcement learning, enhancing algorithmic performance while fostering more interpretable, human-like learning systems. It underscores the potential of cognitive principles to drive more efficient machine learning solutions.
Appendix A Pseudo Code
A.1 Cognitive Belief-Driven Q-Learning (CBDQ) Algorithm
Algorithm 2 Cognitive Belief-Driven Q-Learning Algorithm
[…]
0:  $Q^{CBDQ}(s,a;\phi_{T})$
1:  Initialize $Q(s,a;\phi)$ with random weights $\phi_{0}$;
2:  Initialize replay buffer $\mathcal{B}$ with a fixed length; Initialize Conceptual Experience Organization Framework (CEOF) categories $\{\mathcal{C}_{n}\}^{N}_{n=1}$;
3:  Initialize a $\epsilon$-greedy exploration procedure: Explore($\cdot$)
4:  for each episode do
5:     Get initial state $s_{0}$ from the environment
6:     for each timestep do
7:        Choose action $a_{t}$ using $\epsilon$-greedy: $a_{t}\sim\mathcal{U}(0,1)$
8:        Execute $a_{t}$ to get reward $r(s_{t},a_{t})$, next state $s_{t+1}$
9:        Store $(s_{t},a_{t},r(s_{t},a_{t}),s_{t+1})$ into $\mathcal{B}$
10:        Identify the conceptual category $\mathcal{C}_{i}$ of $s_{t}$, update the count of $a_{t}$ in $\mathcal{C}_{i}$;
11:        Sample $N$ tuples from $\mathcal{B}$ to update $Q$ function:
12:          $y^{i}_{s_{t},a_{t}}=\mathbb{E}_{\mathcal{B}}\left[r(s_{t},a_{t})+\gamma\sum_{a}b_{t}(a\mid s_{t+1})Q(s_{t+1},a;\phi^{-})|s_{t},a_{t}\right]$
13:          The computation of $b_{t}(a\mid s_{t+1})$ in Equation dynamically integrates rewards and subjective beliefs, enabling continuous adaptation based on evolving information.
14:          $Loss=\mathbb{E}_{\mathcal{B}}\left[(y_{s_{t},a_{t}}^{i}-Q(s_{t},a_{t};\phi))^{2}\right]$
15:        Reset after a few updates: $\phi^{-}=\phi$;
@@ A.1 Cognitive Belief-Driven Q-Learning (CBDQ) Algorithm
17:  end for
A.2 Cognitive Belief-Driven Soft Actor-Critic (CBDSAC) Algorithm
Algorithm 3 Cognitive Belief-Driven Soft Actor-Critic
1:  Initialize critic parameters $\phi$, $\bar{\phi}$ and actor parameters $\theta$
2:  Initialize conceptual categories $\{C_{n}\}_{n=1}^{N}$
3:  Initialize category belief parameters $\{\mu_{k},\sigma_{k}^{2}\}_{k=1}^{N}$
4:  for each iteration do
5:     for each environment step do
6:        Identify category $C_{k}$ containing $s_{t}$
7:        Compute $b_{t}(\cdot\mid s_{t})=(1-\beta_{t})\pi_{\theta}(\cdot\mid s_{t})+\beta_{t}\mathcal{P}_{k}(\cdot\mid s_{t})$
8:        Sample $a_{t}\sim b_{t}(\cdot\mid s_{t})$
9:        Transition to $s_{t+1}\sim p(s_{t+1}\mid s_{t},a_{t})$
10:        Store transition in replay buffer: $\mathcal{B}\leftarrow\mathcal{B}\cup\{(s_{t},a_{t},r(s_{t},a_{t}),s_{t+1})\}$
11:        Update category belief parameters:
12:        $\mu_{k}\leftarrow\frac{\sigma_{\text{prior}}^{2}\mu_{\text{obs}}+\sigma_{\text{obs}}^{2}\mu_{\text{prior}}}{\sigma_{\text{prior}}^{2}+\sigma_{\text{obs}}^{2}}$
13:        $\sigma_{k}^{2}\leftarrow\frac{\sigma_{\text{prior}}^{2}\sigma_{\text{obs}}^{2}}{\sigma_{\text{prior}}^{2}+\sigma_{\text{obs}}^{2}}$
14:     end for
15:     for each gradient step do
16:        The critic loss function is minimized as:
$L=\frac{1}{N}\sum_{i=1}^{N}(y_{i}-Q_{\phi}(s_{i},a_{i}))^{2}$
17:        Use parameter updates based on action $a_{i}$, then update the $\theta$ with Equation 19.
18:        Update the temperature coefficient $\alpha$.
19:        $\bar{\phi}\leftarrow\tau\phi+(1-\tau)\bar{\phi}$
@@ A.2 Cognitive Belief-Driven Soft Actor-Critic (CBDSAC) Algorithm
21:  end for
A.3 Cognitive Belief-Driven Proximal Policy Optimization (CBDPPO) Algorithm
@@
Algorithm 4 Cognitive Belief-Driven Proximal Policy Optimization
@@
1:  Initialize policy parameters $\theta_{0}$ and value function parameters $\phi_{0}$
2:  Initialize conceptual categories $\{C_{n}\}_{n=1}^{N}$
3:  for each iteration do
4:     for each environment step do
5:        Collect set of trajectories $D_{k}=\{\tau_{i}\}$ by running $\pi_{k}=\pi(\theta_{k})$
6:        Sample $a_{t}$ and Transition to get $s_{t+1}$
7:        Compute rewards-to-go $r(s_{t},a_{t})$.
8:        Compute advantage estimation $A_{t}$ based on current value function $V_{\phi_{k}}$
9:        Store transition in replay buffer: $\mathcal{B}\leftarrow\mathcal{B}\cup\{(s_{t},a_{t},r(s_{t},a_{t}),s_{t+1},A_{t})\}$
10:     end for
11:     for each gradient step do
12:        Identify category $C_{k}$ containing $s_{t}$
13:        Compute $b_{t}(a\mid s)=(1-\beta_{t})\pi_{\theta}(a\mid s)+\beta_{t}P_{k}(a\mid s)$, based on Equation 20
14:        Update the policy by maximizing the PPO-Clip objective:$\theta_{k+1}=\mathop{argmax}\limits_{\theta}\frac{1}{|D_{k}|T}\sum\limits_{\tau\in D_{k}}\sum\limits_{t=0}^{T}\Big{[}\min\Big{(}\frac{b_{t}(a\mid s)}{\pi_{\theta_{k}}(a\mid s)}A_{t},\text{clip}\left(\frac{b_{t}(a\mid s)}{\pi_{\theta_{k}}(a\mid s)},1-\epsilon,1+\epsilon\right)A_{t}\Big{)}\Big{]},$ based on Equation 21
15:        Fit value function by regression on mean-squared error:$\phi_{k+1}=\mathop{argmax}\limits_{\phi}\frac{1}{|D_{k}|T}\sum\limits_{\tau\in D_{k}}\sum\limits_{t=0}^{T}\Big{(}V_{\phi}(s_{t})-r(s_{t},a_{t})\Big{)}^{2}$
16:     end for
17:  end for
Appendix B Smoothed Bellman Operator
B.1 Lemma
Lemma B.1 (Jensen’s Inequality for Q-values).

### v4

@@ 6 Experiment
Table 2: Average cumulative rewards of HI-PPO, HI-TD3 and baselines across discrete and continuous action environments.
HI-RL for PPO Variants
Method/Environment Atari - AirRaid Atari - Amidar Atari - Asteroids Atari - Centipede Atari - Zaxxon
[…]
HI-RL for TD3 Variants
Method/Environment Box2d - BipedalWalker Mujoco - Ant Mujoco - Swimmer Mujoco - HalfCheetah Mujoco - Walker2d
[…]
Figure 2: Learning curves comparing HI-PPO and PPO (Atari tasks) as well as HI-TD3 and TD3 (Mujoco and Box2D tasks). HI-RL variants demonstrate faster convergence, higher sample efficiency, and reduced variance across diverse environments.
6.1 Comparative Performance of HI-RL and Baselines
To rigorously evaluate the HI-RL framework, we report results across a broad set of benchmark environments spanning both discrete and continuous action spaces (Table 1, Table 2). The tasks range from low-dimensional control (Classic Control, Box2D) to high-dimensional, perceptually rich domains (MetaDrive, MuJoCo), enabling a systematic assessment of generalization and sample efficiency under varying levels of complexity.
Discrete Action Space: As shown in Table 1, HI-DQN (HI-Q) consistently outperforms baselines (DQN, DDQN, Dueling DQN, PPO) across diverse discrete-action tasks. In simple settings such as CartPole, HI-DQN nearly reaches the performance ceiling with lower variance. In more complex tasks like Box2D-CarRacing and MetaDrive, HI-DQN achieves the highest rewards across all sub-tasks, demonstrating robustness and adaptability. Even in intermediate (TOrSX) and highly challenging scenarios (XTOC), HI-DQN maintains clear advantages, highlighting the effectiveness of belief-guided abstraction for stable learning under increasing complexity.
Continuous Action Space: A similar trend is observed in continuous-control benchmarks (Table 1). HI-SAC consistently outperforms SAC, PPO, and DDPG across both medium- and high-dimensional MuJoCo and Box2D tasks. In challenging domains such as Humanoid and HumanoidStandup, HI-SAC achieves substantially higher rewards with improved stability, while in locomotion tasks (HalfCheetah, Walker2d), it converges faster and produces more resilient policies. Overall, these results demonstrate that HI-SAC leverages belief-guided generalization to deliver reliable gains in environments requiring both precise control and long-horizon reasoning.
6.2 Learning Dynamics with Experience-Driven Abstraction
While the previous section demonstrates that HI-RL achieves superior final performance over baseline algorithms in both discrete and continuous action spaces, practical reinforcement learning often places greater emphasis on sample efficiency, training stability, and convergence speed than on post-convergence metrics. These factors are especially critical in resource-constrained or high-risk settings. To this end, we analyze the learning dynamics of HI-PPO vs. PPO and HI-TD3 vs. TD3 on Atari and MuJoCo (Fig. 2, Table 2), illustrating how HI-RL leverages cognitive belief priors for faster exploration and structured abstraction for more stable optimization.
In high-dimensional visual environments such as Atari, HI-PPO consistently improves both convergence speed and final performance. For example, in Amidar, HI-PPO surpasses $2000$ reward at 4040M steps, whereas PPO converges around $\sim 900$. In more challenging tasks such as Asteroids and Centipede, HI-PPO not only learns faster but also exhibits reduced variance, indicating more stable policy updates. The progressive increase of $\beta_{t}$ enables HI-PPO to exploit conceptual priors early on and transition smoothly to task-specific fine-tuning, resulting in efficient and robust learning.
Similarly, in continuous control tasks, HI-TD3 achieves faster convergence, higher rewards, and greater stability compared to TD3. In simpler tasks such as BipedalWalker, HI-TD3 converges more rapidly and attains comparable or better final performance. In more complex locomotion tasks including Ant, Swimmer, HalfCheetah, and Walker2d, HI-TD3 not only reaches higher asymptotic rewards but also produces smoother learning curves with lower variance. By contrast, TD3 often suffers from slower convergence and mid-training stagnation, underscoring the efficiency and robustness advantages of HI-TD3.
@@ 7 Conclusion
We introduce Conceptual Belief-Informed Reinforcement Learning (HI-RL ), a representation-level framework that organizes experiences into conceptual categories and integrates belief-guided fusion into policy learning. Moving beyond buffer replay and static policy libraries, HI-RL establishes a structured memory that supports abstraction, reuse, and generalization. Across Q-learning, PPO, TD3, and SAC, it consistently improves sample efficiency, final returns, and stability in both discrete and continuous domains. By achieving higher returns with fewer interactions and stabilizing updates, HI-RL also reduces computational cost, underscoring its potential for sustainable and resource-efficient training. More broadly, HI-RL illustrates how cognitive principles—conceptual abstraction and belief—can be operationalized to advance reinforcement learning, shifting the field from raw data manipulation toward structured, human-aligned inference. We view this as a step toward an “Era of Experience,” in which intelligence is grounded in the active organization of interaction history rather than rote prediction from data.
Appendix A Appendix
A.1 Conceptual Belief-Informed Twin Delayed Deep Deterministic Policy Gradient (HI-TD3)
TD3 (Twin Delayed Deep Deterministic Policy Gradient)(fujimoto2018addressing), built upon DDPG(silver2014deterministic), mitigates Q-value overestimation and improves stability via twin Q-networks, delayed updates, and target policy smoothing. Here, we focus only on the actor update. Considering a sample batch $D_{t}=(s_{i},a_{i},r_{i},s_{i}^{\prime})$ at time step t, the actor policy update is defined as:
$\mathbb{E}_{s_{i}\sim D_{t}}\big[Q_{\phi_{\min}}(s_{i},\pi_{\theta}(s_{i}))\big]$ (19)
where $Q_{\phi_{\min}}$ takes the smaller value of the two Q-networks, while $\pi_{\theta}$ denotes the policy that generates actions, with $a_{i}=\pi_{\theta}(s_{i})$. Conceptual Belief-Informed TD3 (HI-TD3) applies the HI-RL fusion rule to deterministic policy gradients, refining the concept-based belief prior $b_{t}(k)$ for each conceptual category $C_{k}$ at time step t as:
$\forall(\nabla,k)\in\nabla\times[K]:\quad b_{t}(k)\simeq\nabla_{a}Q_{\phi_{\min}}(s,a)$ (20)
where $b_{t}(k)$, a directional belief, denotes as $\nabla_{a}Q_{\min}(s,a)$, where the smaller Q-value in TD3 is employed to approximate the gradient serving as its representation. The recorded belief direction is updated using an exponential moving average with normalization:
$b_{t}(k)=\frac{(1-\eta)b_{t-1}(k)+\eta\nabla_{a}Q_{\min}(s,a)}{\|(1-\eta)b_{t-1}(k)+\eta\nabla_{a}Q_{\min}(s,a)\|},\quad a\sim\pi_{\theta}(s),\quad s\in C_{k}$ (21)
where $\eta$ is an exponential moving average constant and $b_{t-1}(k)$ denotes the previously stored directional belief.
In the policy update of HI-TD3, we perform belief fusion updates only on the actor network. At each time step t with sampled tuple $(s_{i},a_{i},r_{i},s_{i}^{\prime})$, the integrated directional belief information $B_{t}(k)$ is denotes as:
$B_{t}(k)=c\frac{(1-\beta)\nabla_{a_{i}}Q_{\min}(s_{i},a_{i})+\beta b_{t}(k)}{\|(1-\beta)\nabla_{a_{i}}Q_{\min}(s_{i},a_{i})+\beta b_{t}(k)\|},\quad s_{i}\in C_{k}$ (22)
where c is a constant used to prevent excessive oscillations if $B_{t}(k)$ becomes too large. Differing from previous usage, $\beta$ is determined by directional similarity, computed as a dot product, and serves as the fusion coefficient:
$\beta=\mathrm{clamp}\bigl(\sum_{k}b_{t}(k)\cdot\nabla_{a_{i}}Q_{\phi_{\min}}(s_{i},a_{i}),0,1\bigr)$ (23)
The directional fusion is performed by combining $B_{t}(k)$ as a perturbation with $a_{i}$:
$a_{\rm blend}=\mathrm{clamp}\bigl(a_{i}+B_{t}(k),-1,1\bigr)$ (24)
The actor minimizes to update policy:
$\mathbb{E}_{s_{i}\sim D_{t}}\big[Q_{\phi_{\min}}(s_{i},a_{blend})\big]$ (25)
Thus, HI-TD3 preserves the HI-RL fusion principle through the actor by blending task-driven gradients with conceptual priors, while the critic remains the standard TD3 update for stability. This makes HI-TD3 a deterministic yet framework-consistent instantiation of HI-RL. The pseudocodes are provided in Appendix A.2.4.
A.2 Pseudo Code
A.2.1 Conceptual Belief-Informed Q-Learning (HI-Q) Algorithm
Algorithm 2 Conceptual Belief-Informed Q-learning (HI-Q) Algorithm
1:Initialization: learning rate $\alpha$, discount factor $\gamma$, running steps $T$, episodes $E$, replay buffer $\mathcal{B}$ and a set of K conceptual categories, denoted as $\{\mathcal{C}_{k}\}^{K}_{k=1}$
2:for each episode do
3:  Get initial state $s_{0}$ from the environment
4:  for each timestep t do
5:   Choose a random action $a_{t}$ with probability $\epsilon$ otherwise take $a_{t}=\arg\max_{a}Q(s_{t},a;\theta)$
6:   Execute $a_{t}$ to get reward $r(s_{t},a_{t})$, next state $s_{t+1}$
7:   Store $(s_{t},a_{t},r(s_{t},a_{t}),s_{t+1})$ into $\mathcal{B}$
8:   Identify the conceptual category $\mathcal{C}_{k}$ of $s_{t}$ through Nearest Neighbor
9:   Update the count of $a_{t}$ in $\mathcal{C}_{k}$ (cf. 5);
10:   Sample NN tuples from $\mathcal{B}$ to update $Q$ function:
11:     Extract $b_{t}(a\mid\mathcal{C}_{k}(s_{t}))$ and integrate with rewards to estimate $B_{t}(a\mid s_{t+1})$ (cf.6)
12:     $y^{i}_{s_{t},a_{t}}=\mathbb{E}_{\mathcal{B}}\left[r(s_{t},a_{t})+\gamma\sum_{a}B_{t}(a\mid s_{t+1})Q(s_{t+1},a;\theta^{-})|s_{t},a_{t}\right]$ (cf.8)
13:     $Loss=\mathbb{E}_{\mathcal{B}}\left[(y_{s_{t},a_{t}}^{i}-Q(s_{t},a_{t};\theta))^{2}\right]$
14:   Reset target network after a few updates: align target Q parameters: $\theta^{-}=\theta$;
15:  end for
@@ A.2.1 Conceptual Belief-Informed Q-Learning (HI-Q) Algorithm
A.2.2 Conceptual Belief-Informed Soft Actor-Critic (HI-SAC) Algorithm
Algorithm 3 Conceptual Belief-Informed Soft Actor-Critic
1:Initialize two critic parameters $\phi_{1}$, $\phi_{2}$ and actor parameters $\theta$, Conceptual categories $\{C_{k}\}_{k=1}^{N}$, category belief parameters $b_{t=0}(k)=\{\mu_{k},\sigma_{k}^{2}\}_{k=1}^{N}$
2:for each time step t do
3:  Sample $a_{t}\sim\pi_{\theta}(\cdot\mid s_{t})$
4:  Transition to $s_{t+1}\sim p(s_{t+1}\mid s_{t},a_{t})$
5:  Store transition in replay buffer: $\mathcal{B}\leftarrow\mathcal{B}\cup\{(s_{t},a_{t},r(s_{t},a_{t}),s_{t+1})\}$
6:  for training step do
7:   Sampled $\{s_{i},a_{i},r_{i},s_{i}^{\prime}\}\leftarrow\mathcal{B}$
8:   Identify category $C_{k}$ for $s_{i}$ through Euclidean distance
9:   Compute $B_{t}(si)$ and $B_{t}(si^{\prime})$, $s_{i},s_{i}^{\prime}\in C_{K}$ (cf. 17)
10:   Update category belief parameters $b_{t}(k)$ (cf. 15)
11:  end for
12:  for each gradient step (cf.18) do
13:   Compute target:
$y_{i}=r_{i}+\gamma\,\mathbb{E}_{a^{\prime}_{i}\sim B_{t}(\cdot|s^{\prime}_{i})}\Big[Q_{\min}(s^{\prime}_{i},a^{\prime}_{i})-\alpha\log B_{t}(a^{\prime}_{i}|s^{\prime}_{i})\Big].$
14:   Update critics $(i=1,2)$:
$L_{\text{CBISAC}}^{\text{critic}}(\phi_{i})=\mathbb{E}_{(s_{i},a_{i})\sim D_{t}}\Big[\big(Q_{\phi_{i}}(s_{i},a_{i})-y_{i}\big)^{2}\Big],$
$\phi_{i}\leftarrow\phi_{i}-\eta_{\phi}\nabla_{\phi_{i}}L_{\text{CBISAC}}^{\text{critic}}(\phi_{i}).$
15:   Update actor:
$L_{\text{CBISAC}}^{\text{actor}}(\theta)=\mathbb{E}_{s_{i}\sim D_{t},\,a_{i}\sim B_{t}}\Big[\alpha\log B_{t}(a_{i}|s_{i})-Q_{\min}(s_{i},a_{i})\Big],$
$\theta\leftarrow\theta-\eta_{\theta}\nabla_{\theta}L_{\text{CBISAC}}^{\text{actor}}(\theta).$
16:   Update temperature:
$L(\alpha)=\mathbb{E}_{s_{i},a_{i}\sim B_{t}}\big[-\alpha\,(\log B_{t}(a_{i}|s_{i})+\mathcal{H}_{\text{target}})\big],$
$\alpha\leftarrow\alpha-\eta_{\alpha}\nabla_{\alpha}L(\alpha).$
17:   Soft update target network:
$\bar{\phi}\leftarrow\tau\phi+(1-\tau)\bar{\phi}.$
18:   Update $b_{t}(k)$ (cf.15)
19:  end for
@@ A.2.2 Conceptual Belief-Informed Soft Actor-Critic (HI-SAC) Algorithm
A.2.3 Conceptual Belief-Informed Proximal Policy Optimization (HIPPO) Algorithm
Algorithm 4 Conceptual Belief-Informed Proximal Policy Optimization
@@ A.2.3 Conceptual Belief-Informed Proximal Policy Optimization (HIPPO) Algorithm
1:Initialize policy parameters $\theta$ and value function parameters $\phi$, conceptual categories $\{C_{k}\}_{k=1}^{N}$
2:for each iteration do
3:  for each environment step t do
4:   Collect set of trajectories $D_{k}=\{\tau_{i}\}$ by running $\pi_{k}=\pi(\theta_{k})$
5:   Sample $a_{t}$ and Transition to get $s_{t+1}$
6:   Compute rewards-to-go $r(s_{t},a_{t})$.
7:   Compute advantage estimation $A_{t}$ based on current value function $V_{\phi_{k}}$
8:   Store transition in replay buffer: $\mathcal{B}\leftarrow\mathcal{B}\cup\{(s_{t},a_{t},r(s_{t},a_{t}),s_{t+1},A_{t})\}$
9:  end for
10:  for each gradient step do
11:   Sampled $\{s_{i},a_{i},r_{i},s_{i}^{\prime}\}\leftarrow\mathcal{B}$
12:   Identify category $C_{k}$ for $s_{i}$ through Euclidean distance
13:   Compute $B_{t}(k)=(1-\beta_{t})\pi_{\theta}(a_{i}\mid s_{i})+\beta_{t}b_{t}(k),\quad s_{i}\in C_{k}$ (cf.10)
14:   Update the policy by maximizing the PPO-Clip objective (cf.11):
15:
$\theta_{k+1}\;=\;\arg\max_{\theta}\;\mathbb{E}_{(s_{i},a_{i})\sim\pi_{\theta_{\mathrm{old}}}}\Bigg[\min\!\Bigg(\frac{B_{t}(a_{i}\mid s_{i})}{\pi_{\theta_{\mathrm{old}}}(a_{i}\mid s_{i})}\,A_{i},\;\operatorname{clip}\!\Big(\frac{B_{t}(a_{i}\mid s_{i})}{\pi_{\theta_{\mathrm{old}}}(a_{i}\mid s_{i})},\,1-\epsilon,\,1+\epsilon\Big)A_{i}\Bigg)\Bigg].$
16:   Fit value function by regression on mean-squared error:
17:
$\phi_{k+1}\;=\;\arg\min_{\phi}\;\mathbb{E}_{(s_{i},a_{i})\sim\pi_{\theta_{\mathrm{old}}}}\Big[\big(V_{\phi}(s_{i})-r(s_{i},a_{i})\big)^{2}\Big].$
18:   Update $b_{t}(k)$ (cf.5)
19:  end for
20:end for
A.2.4 Conceptual Belief-Informed Twin Delayed Deep Deterministic (HI-TD3) Algorithm
Algorithm 5 Conceptual Belief-Informed Twin Delayed Deep Deterministic
1:Initialize actor $\pi_{\theta}(s)$, critics $Q_{\phi_{1}}(s,a),Q_{\phi_{2}}(s,a)$, target networks $\theta^{\prime}\leftarrow\theta$, $\phi_{1}^{\prime}\leftarrow\phi_{1}$, $\phi_{2}^{\prime}\leftarrow\phi_{2}$, Conceptual categories $\{C_{k}\}_{k=1}^{N}$, discount factor $\gamma$
2:for each iteration do
3:  for each environment step t do
4:   Sample $a_{t}\;=\;\pi_{\theta}(s_{t})+\epsilon,\quad\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$
5:   Transition to $s_{t+1}\sim p(s_{t+1}\mid s_{t},a_{t})$
6:   Store transition in replay buffer: $\mathcal{B}\leftarrow\mathcal{B}\cup\{(s_{t},a_{t},r(s_{t},a_{t}),s_{t+1})\}$
7:  end for
8:  for each gradient step do
9:   Sample minibatch $D_{t}=\{(s_{i},a_{i},r_{i},s^{\prime}_{i})\}$ from replay buffer
10:   Critic update (TD3)
11:     Compute target action with smoothing: $a_{i}^{\prime}=\pi_{\theta^{\prime}}(s^{\prime}_{i})+\epsilon,\quad\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$
12:     Compute temporal difference target: $y_{i}=r_{i}+\gamma\cdot\min\big(Q_{\phi_{1}^{\prime}}(s^{\prime}_{i},a_{i}^{\prime}),Q_{\phi_{2}^{\prime}}(s^{\prime}_{i},a_{i}^{\prime})\big)$
13:     Update critics by minimizing loss: $L=\frac{1}{N}\sum_{i}\big(y_{i}-Q_{\phi_{j}}(s_{i},a_{i})\big)^{2},\quad j=1,2$
14:   Actor update with fusion
15:     Compute gradient direction: $g_{t}=\nabla_{a}Q_{\min}(s_{i},a_{i})$
16:     Compute fusion coefficient $\beta$ (cf.23)
17:     Integrate belief: $B_{t}=c\cdot\frac{(1-\beta)g_{t}+\beta b_{t}(k)}{\|(1-\beta)g_{t}+\beta b_{t}(k)\|}$ (cf.25)
18:     Blend action: $a_{\mathrm{blend}}=\mathrm{clamp}(\pi_{\theta}(s_{i})+B_{t},-1,1)$(cf.24)
19:     Update actor by minimizing: $J(\theta)=-\frac{1}{N}\sum_{i}Q_{\min}(s_{i},a_{\mathrm{blend}})$(cf.21)
20:   Target network updates
$\phi^{\prime}\leftarrow\tau\phi+(1-\tau)\phi^{\prime},\quad\theta^{\prime}\leftarrow\tau\theta+(1-\tau)\theta^{\prime}$
21:   Update $b_{t}(k)$ (cf.22)
22:  end for
23:end for
A.3 Smoothed Bellman Operator
To reflect cognitive properties of uncertainty-aware decision-making in reinforcement learning, we revise the classical Bellman operator, which updates values deterministically:
$\mathcal{T}Q(s_{t},a_{t})=r_{t}+\gamma\max_{a\in\mathcal{A}}Q_{t}(s_{t+1},a).$ (26)
Here, $\mathcal{T}$ is the classical Bellman operator, $s_{t}\in\mathcal{S}$ denotes the current state, $a_{t}\in\mathcal{A}$ the chosen action, $r_{t}\in\mathbb{R}$ the immediate reward, $\gamma\in(0,1)$ the discount factor, $\mathcal{A}$ the action space, and $Q_{t}(s,a)$ the action–value function at iteration $t$. The $\max_{a\in\mathcal{A}}$ term represents greedy action selection, i.e., propagating value based on the action with the highest estimated return.
$\mathcal{T}_{\text{Smoothed}}Q(s_{t},a_{t})=r_{t}+\gamma\sum_{a\in\mathcal{A}}q_{t}(a\mid s_{t+1})Q_{t}(s_{t+1},a),$ (27)
where $\mathcal{T}_{\text{Smoothed}}$ denotes the Smoothed Bellman Operator, which replaces the hard maximization with a belief-weighted expectation. Here, $q_{t}(a\mid s_{t+1})$ is the action-preference distribution at state $s_{t+1}$, e.g., a softmax distribution over $Q_{t}(s_{t+1},a)$ or the belief-preference distribution in HI-RL. Unlike the deterministic $\max$, this formulation propagates value in a probabilistic, uncertainty-aware manner, balancing task-driven estimates with belief-informed priors.
This smoothing relaxes the deterministic backup, enabling value propagation to account for uncertainty and preference variability. The Smoothed Bellman Operator thus provides a unified, differentiable mechanism for propagating reward uncertainty. In Section 5, we illustrate how Smoothed Bellman Operator integrates with different policy learning paradigms (Q-learning, SAC, PPO). Formal instantiations such as softmax smoothing, clipped interpolation, and Bayesian fusion are provided in next, along with a convergence proof and a Jensen-type inequality.
Lemma A.1 (Jensen’s Inequality for Q-values).